# Spiking Neural Network Architecture Search: A Survey

### v1

@@ III-D Co-Exploring Neural Architecture Search Space and Hardware Design Space
Grid Search HASNAS [22] Systematically evaluates all hyperparameter combinations. Parallelizable and simple to implement.
Ensures the optimum architecture is found.
Computationally intensive.
Impractical for large search spaces.
Random Search SNASNet [113], Liu & Yi [20], Putra & Shafique [21] Selects random combinations of hyperparameters from the search space and evaluates the resulting architecture performances. Explores the search space broadly, finds good configurations faster, and often achieves near-optimal solutions with fewer evaluations.
Simple to implement and parallelizable.
Lacks guided exploration.
Does not adaptively focus on promising regions of the search space.
Evolutionary Zhou et al. [13], AutoSNN [7], GAQ [16], Auto-Spikformer [17], NeuEvo [114], MixedSNN [19], ANAS [77], MSE-NAS [23], DESPINE [27], AutoST [28], SSTNAS [81] Iteratively selects, changes, and reproduces neural networks to optimize a fitness value related to an objective like classification accuracy. Does not require differentiable fitness functions.
Suitable for multi-objective optimization.
Computationally expensive with more parameters.
Reinforcement Learning ANCoEF [84] Rewards positive behaviors and penalizes negative ones to derive efficient strategies toward a given goal.
Balances exploration and exploitation.
Good for complex and high-dimensional search spaces.
Custom reward functions align search with specific goals.
Computationally intensive.
Bayesian Optimization SpikeExplorer [26] Finds the optimum of expensive and poorly understood functions using a probabilistic model-based approach. Requires fewer evaluations.
Capable of finding global optima in complex, multimodal search spaces.
Learns from prior evaluations to focus on promising regions.
Not well-suited for high-dimensional spaces.
Computationally intensive and difficult to parallelize.
Can converge to local optima in noisy settings.
Gradient-Based SpikeDHS [15], Yan et al. [24], LitE-SNN [25], Che et al. [82] Allows for continuous relaxation of the architecture space, enabling gradient-based optimization. Fast exploration, scalable to large spaces.
Efficient for complex architecture searches.
May under-explore the search space.
Susceptible to local optima.
Curriculum Begins with a simple neural network and progressively increases complexity by sampling and evaluating new candidate architectures. Efficient and effective across different data types.
Increases the likelihood of discovering optimal architectures.
Dependent on data quality.
Implementation is complex.
Limited generalization.
Progressive Nijenhuis [14], SpikeNAS [12] Gradually expands the search space, starting with simpler architectures. Parallelizable and less computationally expensive.
Handles multiple resource constraints.
Improved search performance.
Shallow initial architectures may not reflect deeper architecture performance.
Costly for large search spaces.
@@ III-E Neural Network Architecture Search Strategies
Neural architecture search involves exploring a vast design space to identify high-performing configurations for a specific task while minimizing computational resources and time. A successful strategy must evaluate accuracy (often by training on a reduced dataset subset), estimate hardware cost (via real-world measurement or modeling), and select an appropriate search algorithm (e.g., grid search, random search, evolutionary algorithms, reinforcement learning, Bayesian optimization, or gradient-based methods) [93]. Here we briefly describe different popular NAS strategies, with an emphasis on their applicability to SNNAS, as summarized in Table III.
@@ III-E1 Grid Search
Grid search systematically explores every combination of layer counts, types, and hyperparameters to identify an optimal architecture. Although simple and effective for small or constrained search spaces [115], it quickly becomes impractical when the number of configurations grows exponentially. Constrained grid search narrows the range of hyperparameters to control computational cost, but advanced methods (e.g., evolutionary algorithms, reinforcement learning) are typically used when the search space is large. In SNNaS, HASNAS [22] demonstrates how exhaustive methods can be adapted by removing low-impact SNN operations and introducing hardware constraints (e.g., memory, latency, area, energy) prior to the search. This strategy yields an $11\times$ speedup compared to SNASNet [8] on CIFAR10 and CIFAR100 while meeting strict hardware limits and maintaining accuracy, underlining grid search’s viability when the space can be reduced effectively.
@@ III-E2 Random Search
Random search samples configurations from the architecture space at random, forgoing exhaustive exploration and often discovering novel designs missed by more structured methods [116, 117, 118, 119]. Although it does not guarantee quick convergence on optimal solutions, it can yield strong results for SNNs when combined with additional constraints or metrics. For example, Kim et al.[8] introduced a training-free measure of spike activation diversity to identify promising random SNN topologies, resulting in SNASNet. Putra and Shafique[21] adapted this approach to construct memory-efficient networks, showing that larger kernels boost accuracy at the expense of increased memory usage. In another study, Liu et al. [20] randomly sampled 1,000 architectures for the Loihi chip, selected 20 based on gradient-based metrics, then fully trained and evaluated these candidates. Their final architectures achieved competitive accuracy, lower energy consumption, and compact designs compared to state-of-the-art SNNs, highlighting the versatility of random search when guided by well-chosen criteria or hardware-specific goals.
@@ III-E3 Evolutionary Methods
Evolutionary NAS methods (Figure 7) are rooted in early works [120, 121, 122] that mimic biological evolution by iteratively selecting, mutating, and recombining neural network configurations to optimize a fitness function—often accuracy, energy usage, or other objectives [123]. Unlike gradient-driven approaches, they do not require a differentiable objective, making them suitable for multi-objective and discrete design challenges. Genetic Algorithms (GAs) are a common technique that encode networks as strings or graphs, then repeatedly apply crossover and mutation to discover high-performing architectures. Historically, these methods have optimized weights (e.g., NEAT [72]), and more recent works extend them to structural and hyperparameter searches [124, 125, 126], sometimes using illumination algorithms like MAP-Elites [127] for diverse solutions.
Despite high computational demands (e.g., AmoebaNet-A [128] required 3,150 GPU days, evolutionary algorithms can be accelerated through smaller training sets, prediction models, Lamarckian strategies [129, 130], or partial weight inheritance. Notable ANN HW-NAS efforts include TinyNAS [131], Once-for-all [132], HAT [133], NASCaps [134], and HURRICANE [135]. For SNNs, evolutionary approaches are similarly powerful for tasks requiring hardware, memory, and energy optimizations, as demonstrated by GAQ-SNN [16], DESPINE [27], MixedSNN [19], MSE-NAS [23], Auto-Spikformer [17], NeuEvo [18], and ANAS [77], among others. Their parallelizable, global search can reveal novel SNN designs in scenarios where gradient signals are limited, although balancing computation, efficient encoding, and hyperparameter tuning remains a key challenge.
Flow diagram of an evolutionary algorithm for neural architecture search. Three large boxes contain sets of architectures, shown as small colored node-and-edge graphs, and are connected in a closed loop. The top-left box, labeled ’Population,’ is the initial pool of candidate architectures. The word ‘Initialization’ with an arrow points to this box to indicate the population is randomly or heuristically initialized. An arrow labeled ’Parent Selection’ leads from ’Population’ to the box on the right labeled ’Parents,’ showing that a subset of candidates is chosen as parent architectures. An arrow labeled ’Crossover and Mutation’ leads from ’Parents’ to the bottom-center box labeled ’Children,’ indicating that genetic operators are applied to the parents to produce offspring. Finally, an arrow labeled ’Evaluation, Selection, and Update’ leads from ’Children’ back to ’Population,’ indicating that the child architectures are evaluated for performance, the best are selected, and the population is updated for the next generation. Overall workflow: initialize a population, select parents, apply crossover and mutation to create children, evaluate and select the best, update the population, and repeat.
@@ III-E4 Reinforcement Learning Methods
Reinforcement learning (RL) methods, illustrated in Figure 8, frame NAS as a Markov Decision Process [137], where an agent explores network architectures (or hardware designs) by taking actions that yield rewards based on accuracy or efficiency. Early work by Zoph and Le [78] introduced RL-based controllers for NAS, but required extensive computational resources; subsequent techniques such as weight sharing [76] and learning curve prediction [138] reduced this overhead. RL is appealing for hardware-oriented NAS because controllers can dynamically adjust to multi-objective goals (e.g., latency, energy) [94, 139, 140, 90, 111].
@@
Despite these advances, RL in SNNaS remains relatively unexplored. Challenges include high computational demands, large action spaces, and potential local optima [141, 128, 142]. RL controllers can be sample-inefficient and may require extensive tuning to generalize across tasks. Future directions involve more robust environment designs, refined reward functions, and possible hybridization with evolutionary or Bayesian methods to mitigate sample intensity and improve scalability.
Reinforcement-learning loop for neural architecture search. At the bottom center is a box labeled ‘The Controller (RNN)’ showing a small recurrent neural network. From the controller, a right-pointing arrow labeled ‘Action (Sample Architecture)’ goes to the top-center box labeled ‘Environment (Neural Architecture),’ which contains a small directed graph representing a candidate network. From the environment, a left-pointing arrow labeled ‘Reward (Accuracy on Validation Set)’ returns to the controller. The two arrows form a closed loop: the controller samples an architecture (action), the environment trains/evaluates it, and the resulting validation accuracy is fed back as a reward to update the controller’s policy.
Figure 8: Illustration of reinforcement learning for NAS based on [78].
@@ III-E5 Bayesian Optimization
Bayesian Optimization (BO) adaptively searches expensive, complex objective functions by building a probabilistic surrogate model (e.g., a Gaussian process) and balancing exploration and exploitation via an acquisition function [143]. In SNNaS, BO typically encodes architectural elements (e.g., neuron types, connectivity) and performance metrics (e.g., accuracy, sparsity) into a surrogate model that guides the next candidate selection, thereby minimizing costly simulation and training runs. The approach iteratively updates the surrogate model with newly evaluated architectures, refining predictions and uncertainties; careful tuning of hyperparameters, surrogate design, and acquisition functions remains essential. Despite potential computational overhead and risks of local optima, BO’s data-efficient search can yield high-quality solutions, as shown by SpikeExplorer [26], which uses BO to co-optimize power, latency, area, and accuracy on FPGAs. This enables efficient exploration across multiple constraints, ultimately outperforming several accelerator baselines while preserving accuracy.
@@ III-E6 Gradient-Based Methods
Gradient-based NAS methods optimize architecture parameters via backpropagation through a continuous relaxation of discrete choices [144]. Early examples include DARTS [144], SNAS [113], ProxylessNAS [145], and FBNet [146], which reduced the memory overhead and search time by restricting continuous activation paths or using stochastic gradient updates. However, some studies, such as Bingham et al.[147], report that gradient descent may stagnate on a single architecture over many iterations, failing to exploit the broader search space as effectively as random or evolutionary methods. Despite these challenges, gradient-based strategies remain popular due to their efficiency once the continuous relaxation is set up, and they are frequently combined with techniques like inference latency constraints, as in HTAS[148].
For SNNs, gradient-based methods have spawned frameworks like SpikeDHS [15], which adapts DARTS for spiking neurons by addressing surrogate gradient noise and maintaining multiplication-free inference. Che et al.[82] further enhance SpikeDHS with differentiable surrogate gradient search (DGS) and temporal parameter search (TPS) to refine neuron dynamics, and these modules can be applied individually or jointly for added flexibility. LitE-SNN[25] extends SpikeDHS to incorporate spatial and temporal compression, and Yan et al. [24] propose a branchless supernet that encodes architectures with continuous parameters for gradient-based multi-objective optimization. Collectively, these approaches have achieved state-of-the-art performance across various datasets by carefully balancing the high computational demands of differentiable searches with specialized SNN design considerations.
III-E7 Progressive Neural Architecture Search
Progressive Neural Architecture Search (PNAS)[149] incrementally expands a base network by adding layers or connections, focusing computational effort on promising partial architectures. This “search by growing” strategy drastically reduces the exploration of unproductive designs and often delivers results significantly faster. PNAS can be five times faster than earlier RL-based methods[150] and calculate the results eight times faster overall [149]. Despite its efficiency and strong performance on benchmarks like CIFAR-10 and ImageNet, PNAS can prematurely prune architectures and incur high training costs for performance prediction [151]. In SNNaS, adaptations include the Sequential Greedy Architecture Search (SGAS) used by Nijenhuis [14], which did not surpass baseline SNN results, and SpikeNAS [12], a memory-aware PNAS variant that progressively refines cell-based SNN designs by pruning memory-intensive operations. These studies illustrate both the potential and the challenges of applying PNAS to spike-based architectures.
III-E8 Curriculum Neural Architecture Search
CNAS [151] employs a staged, curriculum-inspired process to navigate complex NAS landscapes. It begins with a simple base network, then gradually adds operations or layers via mutation or component combination. Newly generated architectures are briefly trained and compared to the base; stronger performers supersede earlier models. By focusing resources on promising partial solutions, CNAS can achieve better convergence within limited budgets, splitting NAS into manageable subproblems that exploit learned knowledge from previous stages.
III-E9 Hybrid Methods
@@
Hybrid approaches blend multiple NAS strategies to combine their benefits, such as robust exploration from one method and efficient exploitation from another, thereby speeding searches and conserving resources. Examples include evolutionary algorithms coupled with reinforcement learning [152, 153, 154] and random search paired with local grid search. Although these hybrids can produce more balanced results and handle diverse tasks, they also introduce increased complexity and hyperparameter tuning. Careful design is crucial to avoid overlapping strengths and redundant computations, but when done well, hybrid NAS can outperform single-method solutions.

### v2

@@ III-D3 Asynchronous and Event-Driven Hardware Constraints
Grid Search NeuroNAS [111], ANAS [179], [4] Systematically evaluates all hyperparameter combinations. Parallelizable and simple to implement. Ensures the optimum architecture is found. Computationally intensive. Impractical for large search spaces.
Random Search SNASNet [68], Liu & Yi [84], Spikernel [112], STMixer [42], ESL-SNNs [133], ANAS [179]1, ESNN [170]1, TNAS-ER [83]1, Auto-Spikformer [30]2, AutoSNN [97]2 Selects random combinations of hyperparameters from the search space and evaluates the resulting architecture performances. Explores the search space broadly, finds good configurations faster, and often achieves near-optimal solutions with fewer evaluations. Simple to implement and parallelizable. Lacks guided exploration. Does not adaptively focus on promising regions of the search space.
Evolutionary AutoSNN [97], GAQ [99], Auto-Spikformer [30], NeuEvo [134], MixedSNN [167], ANAS [179], MSE-NAS [104], DESPINE [5], AutoST [160], SSTNAS [74], EONS [128], ABC-Arc [182], EMO-SNAS [138], EB-NAS [105], PESNN [159], SPENSER [25], BCNAS-SNN [40], MONAS-ESNN [125], GP-Assisted CMA-ES [186], Membrane-NAS [78], Spiking WANN [11] Iteratively selects, changes, and reproduces neural networks to optimize a fitness value related to an objective like classification accuracy. Does not require differentiable fitness functions. Suitable for multi-objective optimization. Computationally expensive with more parameters.
Deep Reinforcement Learning (DRL) ANCoEF [180], SPEAR [165], GAQ [99] Rewards positive behaviors and penalizes negative ones to derive efficient strategies toward a given goal. Balances exploration and exploitation. Scales to complex and high-dimensional search spaces. Flexible reward design enables custom multi-objective optimization. Computationally expensive and sample-inefficient. Compounded latency with SNNs. Training instability and tuning sensitivity. Additional controller infrastructure required. Increased latency in SNN settings.
Bayesian Optimization (BO) SpikeExplorer [102], ASNPC [157], GP-Assisted CMA-ES [186], Benmeziane et al. [16], deepHyper [173] Finds the optimum of expensive and poorly understood functions using a probabilistic model-based approach. Requires fewer evaluations. Capable of finding global optima in complex, multi-modal search spaces. Learns from prior evaluations to focus on promising regions. Not well-suited for high-dimensional spaces. Computationally intensive and difficult to parallelize. Can converge to local optima in noisy settings.
Gradient-Based SpikeDHS [29, 31], ESNN [170], LitE-SNN [82], MA-DARTS [90], ST-DANO [178], SpikeDHS [29, 31] Allows for continuous relaxation of the architecture space, enabling gradient-based optimization. Fast exploration, scalable to large spaces. Efficient for complex architecture searches. May under-explore the search space. Susceptible to local optima.
  • 1
    Implemented in comparative baseline.
  • 2
    Implemented in ablation studies.
@@ III-E Neural Network Architecture Search Strategies
NAS involves exploring a vast design space to identify high-performing configurations for a specific task while minimizing computational resources and time. A successful strategy must evaluate performance (often by training on a reduced dataset subset), estimate hardware cost (via real-world measurement or modeling), and select an appropriate search algorithm (e.g., grid search, random search, evolutionary algorithms, deep reinforcement learning, Bayesian optimization, or gradient-based methods) [15]. Here we briefly describe different popular NAS strategies, with an emphasis on their applicability to SNNaS, as summarized in Table III.
@@ III-E1 Grid Search
Grid search systematically explores all hyperparameter combinations but becomes impractical as search spaces grow exponentially [75]. NeuroNAS [111], an SNNaS framework, adapts the grid search by pruning low-impact SNN operations and incorporating hardware constraints directly, achieving $8.9\times$ speedup over SNASNet [68] while maintaining accuracy on CIFAR and TinyImageNet.
@@ III-E2 Random Search
Random search samples configurations at random, often discovering novel designs missed by structured methods [17, 176, 71, 116]. SNASNet [68] uses initialization-time, training-free (no weight updates) spike activation diversity metrics to identify promising random SNN topologies. Liu et al. [84] randomly sampled 1,000 architectures for Loihi, and trained the top 20 according to gradient metrics to achieve competitive accuracy with lower energy than state-of-the-art SNNs.
@@ III-E3 Evolutionary Methods
Evolutionary NAS methods (Figure 4) are rooted in early works [95, 51, 150] that mimic biological evolution by iteratively selecting, mutating, and recombining neural network configurations to optimize a fitness function—often accuracy, energy usage, or other objectives [86, 128]. Unlike gradient-driven approaches, they do not require a differentiable objective, making them suitable for multi-objective and discrete design challenges [128]. Genetic Algorithms (GAs) are a common technique that encode networks as strings or graphs, then repeatedly apply crossover and mutation to discover high-performing architectures [128, 5, 25]. Historically, these methods have optimized weights (e.g., NeuroEvolution of Augmenting Topologies (NEAT) [140]), and more recent works extend them to structural and hyperparameter searches [135, 175, 120], sometimes using illumination algorithms like Multi-dimensional Archive of Phenotypic Elites (MAP-Elites) [96] for diverse solutions.
Despite high computational demands (e.g., AmoebaNet-A [119] in ANN NAS required 3,150 GPU days), evolutionary algorithms can be accelerated through smaller training sets, prediction models, Lamarckian strategies [110, 45], or partial weight inheritance. Notable ANN HW-NAS efforts include TinyNAS [76], Once-for-all [27], Hardware-Aware Transformers (HAT) [155], NAS for convolutional capsule networks (NASCaps) [91], and HURRICANE [181]. For SNNs, evolutionary approaches are similarly powerful for tasks requiring hardware, memory, and energy optimizations, as demonstrated by EONS [128], GAQ-SNN [99], DESPINE [5], MixedSNN [167], MSE-NAS [104], Auto-Spikformer [30], NeuEvo [132], and ANAS [179], among others. Their parallelizable, global search can reveal novel SNN designs in scenarios where gradient signals are limited, although balancing computation, efficient encoding, and hyperparameter tuning remains a key challenge.
@@ III-E4 Reinforcement Learning Methods
Reinforcement learning (RL) methods, illustrated in Figure 5, frame NAS as a Markov Decision Process [144], where an agent explores network architectures (or hardware designs) by taking actions that yield rewards based on accuracy or efficiency. In practice, most RL-based NAS methods effectively employ deep reinforcement learning (DRL), where policies are parameterized by neural networks to manage high-dimensional, discrete design spaces. Early work by Zoph and Le [187] introduced DRL-based controllers for NAS but required extensive computational resources; subsequent techniques, such as weight sharing [109] and learning curve prediction [13] reduced this overhead. DRL is appealing for hardware-oriented NAS because controllers can dynamically adjust to multi-objective goals (e.g., latency, energy) [184, 146, 88, 2, 64].
@@
Despite these advances, DRL-based SNNaS methods are relatively unexplored. Challenges include high computational demands, large action spaces, potential local optima, and dynamic SynOps [86, 121, 138, 165]. DRL controllers can be sample-inefficient and may require extensive tuning to generalize across tasks [165]. Future directions involve more robust designs, refined reward functions, and possible hybridization with evolutionary or BO methods to mitigate sample intensity and improve scalability.
Figure 5: Illustration of RL for NAS based on [187].
@@ III-E5 Bayesian Optimization
BO adaptively searches expensive, complex objective functions by building a probabilistic surrogate model (e.g., a Gaussian process) and balancing exploration and exploitation via an acquisition function [130]. In SNNaS, BO typically encodes architectural elements (e.g., neuron types, connectivity) and performance metrics (e.g., accuracy, sparsity) into a surrogate model that guides the next candidate selection, thereby minimizing costly simulation and training runs. The approach iteratively updates the surrogate model with newly evaluated architectures, refining predictions and uncertainties; careful tuning of hyperparameters, surrogate design, and acquisition functions remains essential. Despite potential computational overhead and risks of local optima, BO’s data-efficient search can yield high-quality solutions, as shown by SpikeExplorer [102], which uses BO to co-optimize power, latency, area, and accuracy on FPGAs. This enables efficient exploration across multiple constraints, ultimately outperforming several accelerator baselines while preserving accuracy.
@@ III-E6 Gradient-Based Methods
Gradient-based NAS methods optimize architecture parameters via backpropagation by continuously relaxing discrete choices [81]. Early examples include Differentiable Architecture Search (DARTS) [81], Stochastic NAS (SNAS) [166], ProxylessNAS [28], and FBNet (a hardware-aware efficient ConvNet design via differentiable NAS) [163], which reduced the memory overhead and search time by restricting continuous activation paths or using stochastic gradient updates. However, some studies, such as Bingham et al. [21], report that gradient descent may stagnate on a single architecture over many iterations, failing to effectively exploit the broader search space as random or evolutionary methods do. Despite these challenges, gradient-based strategies remain popular due to their efficiency once continuous relaxation is established, and they are frequently combined with techniques such as inference latency constraints, as in HTAS [65].
Gradient-based methods are common in SNNaS. SpikeDHS [29] adapts the DARTS [81] framework for SNNs by implementing spike-based computation at cell and layer levels, ensuring multiplication-free inference. It uses Differentiable Surrogate Gradient Search (DGS) to optimize local gradient approximations and prevent the network from getting stuck in local minima. The approach is further extended with Temporal Parameter Search (TPS) and Hybrid Network Search (HNS) to refine neuron dynamics and balance efficiency. LitE-SNN [82] adds spatial and temporal compression to SpikeDHS, and Yan et al. [170] propose a branchless supernet, ESNN, that encodes architectures with continuous parameters for gradient-based multi-objective optimization. MA-DARTS [90], a prime example of Gradient-Based SNNaS, uses bi-level optimization to simultaneously update architecture parameters ($\alpha$) on validation data and network weights (w) on training data. It integrates a multi-dimensional attention mechanism into the node concatenation step, enabling the model to prioritize critical temporal, channel, and spatial information while ignoring less relevant data. This addition significantly improves classification accuracy with only a negligible increase in network size and complexity. Collectively, these approaches have achieved state-of-the-art performance across various datasets by carefully balancing the high computational demands of differentiable searches with specialized SNN design considerations.
III-E7 Hybrid Methods
@@
Hybrid approaches blend multiple NAS strategies to combine their benefits, such as robust exploration from one method and efficient exploitation from another, thereby speeding searches and conserving resources. Examples include evolutionary algorithms coupled with DRL [35, 93] and random search paired with local grid search. Although these hybrids can produce more balanced results and handle diverse tasks, they also introduce increased complexity and hyperparameter tuning. Careful design is crucial to avoid overlapping strengths and redundant computations, but when done well, hybrid NAS can outperform single-method solutions.
III-E8 Search Space Primitives and Inductive Biases
Architectural primitives available within an SNNaS search space act as inductive biases that strongly shape the quality and efficiency of discovered SNNs. Common choices include cell-based DAGs with searchable operations [82, 3, 90, 111], convolutional kernels that balance representational capacity against memory cost [111, 113], and pooling strategies that favor max-pooling to preserve sparsity and binary spike behavior [100, 97, 102]. Unlike ANN NAS, many SNNaS frameworks further expose temporal structure through backward or recurrent connections [68, 125, 153], and searchable neuron models and internal parameters such as thresholds, decay rates, and time constants [167, 31, 5].
In spiking transformer models, attention mechanisms have emerged as an additional inductive bias. MA-DARTS [90] integrates multi-dimensional attention (channel, spatial, and temporal) directly into differentiable search, enabling attention-weighted node aggregation during optimization. Transformer-based approaches such as Auto-Spikformer [30] treat multi-head spiking self-attention as a searchable primitive, while STMixer [42] replaces self-attention with hardware-friendly multi-head token mixing to better integrate temporal and spatial features. By contrast, NAS-TinyML [34] applies channel-wise attention only to guide macro-level decisions rather than searching attention structure. These methods highlight attention—particularly temporal attention—as an emerging bias for modeling spike-timing variability within SNN search spaces.
Additional inductive biases are imposed through biologically inspired modularity [105, 104], explicit efficiency objectives based on spike counts or synaptic operations [97, 165, 125, 138], and hardware-aware constraints that prune unsupported operations [84, 111]. Many frameworks further bias search using training-free proxies that favor diverse initial activations [125, 68, 74], layer-wise alignment of neuron dynamics with temporal dependency requirements [125, 5], or spatial embedding of neurons that constrains connectivity based on physical or biological distance [104, 153, 132].